How many grams of aluminum oxide will be formed upon the complete reaction of 31.6 grams of aluminum with excess iron(III) oxide?


Fe2O3 (s) +2Al (s) -------> Al2O3 (s) +2Fe (s)


..............grams aluminum oxide


please with steps

1
Expert's answer
2021-03-03T06:29:06-0500

Solution:

n(Al) = 31.6/27 = 1.17 mole

n(Al2O3) = 0.585 mole

m(Al2O3) = 0.585×102 = 59.7 g